in a past​ election, the voter turnout was 62​%. In a​ survey, 806 subjects were asked if they voted in the election. a. Find the mean and standard deviation for the numbers of voters in groups of 806. b. In the survey of 806 ​people, 498 said that they voted in the election. Is this result consistent with the​ turnout, or is this result unlikely to occur with a turnout of 62​%? Why or why​ not? c. Based on these​ results, does it appear that accurate voting results can be obtained by asking voters how they​ acted? a. mu μ equals = nothing ​(Round to one decimal place as​ needed.) sigma σ equals = nothing ​(Round to one decimal place as​ needed.)
Answer #1

Solution:

(a) μ = 806(0.62)
= 499.72
σ = √499.72(0.38)
= 13.78

(b) μ + 2σ
= 499.72 + 2(13.78)
= 499.72 + 27.56
= 527.28
498 voters are consistent with 62% because 498 < 527.28

(c) No